OpenStudy (anonymous):

f(x) = 3x + 1; g(x) = 5x – 1 Find f/g

OpenStudy (anonymous):

f/g=3x + 1/5x – 1

OpenStudy (anonymous):

domain?

OpenStudy (anonymous):

all x don't make 5x-1=0 so all x exept x=1/5